# FineScope: SAE-guided Data Selection Enables Domain-Specific LLM Pruning & Fine-Tuning

### v2

@@ Appendix B Seed Selection:
We visualize the seed selection process using cluster analysis, as shown in Figure 6. The interpretation of features varies across initial, middle, and final layers, reflecting how representations evolve within the model. The results suggest that our curated dataset is alinged with the clustered virtual domain target domains across different layer representations. Lower layers tend to capture broad, generalized features, often encoding syntactic structures or common linguistic patterns. In contrast, deeper layers focus on increasingly abstract and domain-specific attributes, leading to more compact and semantically meaningful clusters.
@@
This progressive refinement suggests that domain-relevant information emerges more distinctly in later layers, where feature representations become more specialized. Our method leverages this hierarchical structure, selecting seeds from layers that balance generalization and domain specificity. By utilizing SAE-based representations instead of raw embeddings, we ensure that the seed selection process is more interpretable and aligned with high-level domain knowledge, rather than being influenced by superficial token-level similarities. This reinforces the effectiveness of our approach in identifying the most informative seed samples for dataset curation.
@@
B.1 Example Seeds
Figures 78, and  9 present the user-defined seed samples for STEM, Social sciences and Humanities respectively. Based on these seeds, we extracted domain-specific data points from the larger dataset.

### v3

@@ 14.1 Seed Selection
In this section, we include a qualitative analysis to make the seed selection mechanism more transparent. We visualize the seed selection process using cluster analysis, as shown in Figure 9. The interpretation of features varies across initial, middle, and final layers, reflecting how representations evolve within the model. The results suggest that our curated dataset is alinged with the clustered virtual domain target domains across different layer representations. Lower layers tend to capture broad, generalized features, often encoding syntactic structures or common linguistic patterns. In contrast, deeper layers focus on increasingly abstract and domain-specific attributes, leading to more compact and semantically meaningful clusters.
@@
This progressive refinement suggests that domain-relevant information emerges more distinctly in later layers, where feature representations become more specialized. Our method leverages this hierarchical structure, selecting seeds from layers that balance generalization and domain specificity. By utilizing SAE-based representations instead of raw embeddings, we ensure that the seed selection process is more interpretable and aligned with high-level domain knowledge, rather than being influenced by superficial token-level similarities. This suggests that SAE-guided representations yield more semantically aligned and thus more reliable retrieval than raw embedding similarity.
@@
14.2 Example Seeds
For reproducibility and to clarify the user input required by FineScope, we provide representative seed prompts for each domain; these seeds are the primary human-specified inputs to the pipeline. Figures 1011, and  12 present the user-defined seed samples for STEM, Social sciences and Humanities respectively. Based on these seeds, we extracted domain-specific data points from the larger dataset.